# Segregation in Nuclear Stellar Clusters: Rates and Mass Distributions of TDEs, QPEs, Plunges, and EMRIs

### v1

@@ Abstract
Supermassive black holes at the centers of galaxies occasionally disrupt stars or consume stellar-mass black holes that wander too close, producing observable electromagnetic or gravitational wave signals. We examine how mass segregation impacts the rates and distributions of such events. Assuming a relaxed stellar cluster, composed of stars and stellar-mass black holes, we show that the tidal disruption rate of massive stars ($m\gtrsim{M_{\odot}}$) is enhanced relative to their abundance in the stellar population. For stars up to $m\approx 3{M_{\odot}}$, this enhancement is roughly $m/M_{\odot}$ and it is driven by segregation within the sphere of influence. Stars with masses $m\gtrsim 3{M_{\odot}}$, if relaxed, are predominantly scattered by more massive stellar-mass black holes, leading to a constant enhancement factor of $\approx 9$, independent of mass. This aligns with observational evidence suggesting an over-representation of massive stars in tidal disruption events. For stellar-mass black holes, we predict an enhancement factor scaling as $m_{\bullet}^{1/2}$ for plunges and $m_{\bullet}^{3/2}$ for extreme-mass-ratio inspirals (EMRIs). The power of one-half in both cases reflects the shorter relaxation times of heavier black holes, allowing them to segregate into the sphere of influence from greater distances, thereby increasing their abundance. The additional power in the EMRIs’ rate arises from the tendency of heavier black holes to circularize and sink inward more efficiently. Finally, we estimate the rate of main sequence star inspirals and find that it favors low-mass stars ($m\lesssim{M_{\odot}}$). This seems compatible with the observationally estimated rate of quasi-periodic eruptions.
@@ 1 Introduction
The centers of galaxies harbor supermassive black holes (SMBHs) and their surrounding nuclear stellar clusters. These dense environments give rise to various electromagnetic and gravitational wave (GW) transients, such as tidal disruption events (TDEs; Rees, 1988; Gezari, 2021), quasi-periodic eruptions (QPEs; Miniutti et al., 2019; Arcodia et al., 2021), merging stellar-mass black hole (BH) binaries (Mapelli, 2021; Arca Sedda et al., 2023), and extreme-mass-ratio inspirals (EMRIs; Amaro-Seoane, 2018).
The rates and characteristics of these transients depend on the distribution of stars and BHs around the SMBH, which has been extensively studied over the past half a century, since the pioneering works of Peebles (1972) and Bahcall & Wolf (1976).
Under the assumptions of spatial spherical symmetry, isotropic velocities, and weak two-body scattering dynamics, Bahcall & Wolf (1976) derived a zero-flux steady-state solution for a single-mass population. In this case, the phase-space distribution is given by $f(E)\propto E^{p}$ with $p=1/4$, which corresponds to a number density $n(r)\propto r^{-\eta}$, with $\eta=3/2+p=7/4$ (hereafter, BW profile). This solution satisfies a vanishing particle flux and a constant energy flux (Rom et al., 2023). In a following paper, Bahcall & Wolf (1977) generalized their calculation for multi-mass groups. Assuming that the most massive objects, $m_{\max}$, are the most abundant, a zero-flux solution is satisfied when the massive group follows the single-mass BW profile, with $p_{\max}\approx 1/4$, while lighter objects, with mass $m_{i}<m_{\max}$, obtain shallower profiles (Bahcall & Wolf, 1977), satisfying
$\frac{p_{i}}{p_{\max}}\approx\frac{m_{i}}{m_{\max}}.$ (1)
However, in realistic nuclear stellar clusters, low-mass stars are much more abundant than massive stars or compact objects, hence they dominate the scattering and the massive objects sink toward the center of the cluster, due to dynamical friction, as encapsulated by Alexander & Hopman (2009) and Keshet et al. (2009) mass-segregated distributions. Recently, Linial & Sari (2022) revisited the impact of segregation in nuclear stellar clusters and derived a zero-flux, steady-state solution, taking into account the dominance of different mass groups at different energy bins. Here, we follow this notion to study the rate of various observable transients in galactic centers.
A star is tidally disrupted if it passes too close to the SMBH, namely closer than its tidal radius, where the SMBH tidal force and the star’s self-gravity are comparable (Hills, 1975),
@@
where $m$ and $R_{\star}$ are the mass and the radius of the star, respectively, and $R_{s}=2GM_{BH}/c^{2}$ is the SMBH Schwarzschild radius. We normalize $M_{BH}$ to the mass of Sgr-A*, $M_{\rm MW}=4\times 10^{6}$ (Ghez et al., 2008; Gillessen et al., 2009), and use the main-sequence mass-radius relation, $R_{\star}\propto m^{0.8}$.
As a consequence of such a close encounter, the star is torn apart. Roughly half of its mass is been ejected while the rest remains bound to the SMBH, form an accretion disk and powers a distinctive luminous flare (Rees, 1988). The TDEs rate, $10^{-4}-10^{-5}\ \rm yr^{-1}$ (Magorrian & Tremaine, 1999; Wang & Merritt, 2004; Holoien et al., 2015; Stone & Metzger, 2016; Kochanek, 2016; van Velzen, 2018; Stone et al., 2020; Bortolas et al., 2023; Yao et al., 2023), is determined by the replenishment, mainly via two-body scatterings, of stars into highly eccentric orbits with periapsis $r_{p}\lesssim R_{t}$.
@@
Observationally, about $100$ TDEs have been discovered (Gezari, 2021; Hammerstein et al., 2023; Yao et al., 2023), primarily in the past decade, through wide field surveys, such as the All-Sky Automated Survey for Supernovae (ASAS-SN; Jayasinghe et al., 2018) and the Zwicky Transient Facility (ZTF; Bellm et al., 2019). The number of detected TDEs will significantly increase with the upcoming Vera Rubin Observatory Legacy Survey of Space and Time (LSST; Ivezic et al., 2019), which is expected to detect $\sim 10^{3}$ TDEs per year (van Velzen et al., 2011; Bricman & Gomboc, 2020).
The population of disrupted stars is likely dominated by low-mass stars, with masses around $m\lesssim 1$, as they are more common and have longer lifespans than more massive stars (Kochanek, 2016). However, there are observational indications of disruptions of more massive stars, with masses up to a few solar masses (e.g., Mockler et al., 2022; Hinkle et al., 2024; Wiseman et al., 2024).
@@
Additionally, the recent discoveries of QPEs (Miniutti et al., 2019; Arcodia et al., 2021) uncovered a new class of transients from centers of galaxies, characterized by repeating x-ray emissions over periods of hours. The growing number of observed QPEs reveals the complexity and diversity of phenomena associated with these events (Arcodia et al., 2022; Miniutti et al., 2023; Miniutti, G. et al., 2023; Arcodia et al., 2024a; Chakraborty et al., 2024). Several formation models have been proposed, including mass transfer from white dwarfs (King, 2022) or stars (Metzger et al., 2022; Lu & Quataert, 2023; Linial & Sari, 2023), star-disk interactions (Xian et al., 2021; Franchini et al., 2023; Linial & Metzger, 2023), and disk instabilities (Raj & Nixon, 2021; Pan et al., 2022; Kaur et al., 2023). Several evidences supporting the TDE–QPE connection, as suggested by Linial & Metzger (2023), have been observed (Chakraborty et al., 2021; Quintin et al., 2023; Bykov et al., 2024). Most recently, Nicholl et al. (2024) provided a direct link between these phenomena by detecting QPEs following a known TDE.
@@
In parallel, BHs accumulate in galactic centers and migrate inward due to dynamical friction, eventually merging with the central SMBH. These mergers produce GW signals in the mHz band, detectable by next-generation, space-based GW observatories, such as LISA (Amaro-Seoane et al., 2017, 2023) and TianQin (Luo et al., 2016). For these mergers, the critical distance from the SMBH is $r_{p}=4R_{s}$ (rather than the tidal radius associated with the disruption of stars), corresponding to the angular momentum of the mostly bound orbit (Hopman & Alexander, 2005). BHs that reach a smaller periapsis will rapidly plunge into the SMBH. In contrast, BHs on highly eccentric orbits with larger periapsis distance can dissipate energy via GW emission, circularize, and slowly descend toward the SMBH. Such orbital evolution is known as EMRI (Hopman & Alexander, 2005; Amaro-Seoane, 2018).
@@
We highlight the impact of mass segregation on the rates and mass distribution of the different galacto-centric transients. For example, if the disrupted stars of all masses were to originate from roughly the radius of influence, their relative abundances would correspond to their prevalence in the stellar population. However, if mass segregation occurs, massive objects occupy more tightly bound orbits. The impact of such segregation is widely studied in the context of BHs and its consequent enhancement of EMRI rates (e.g., Hopman & Alexander, 2006; Amaro-Seoane & Preto, 2011; Aharon & Perets, 2016; Raveh & Perets, 2021; Broggi et al., 2022; Rom et al., 2024). Analogously, mass segregation could increase the disruption rate of massive stars. But, unlike BHs, the relevance of segregation for massive stars is uncertain due to their short lifetimes.
@@
Nonetheless, massive stars are found in the center of our galaxy, despite the local relaxation time being longer than their expected ages. This is the known “paradox of youth” (Ghez et al., 2003; Lu et al., 2006; Genzel et al., 2010). Unfortunately, the mechanisms that enables massive stars to reside deep within the sphere of influence in our own galactic center are unclear. Various mechanisms, accounting for the the massive stars’ unexpected presence, were suggested; from in situ formation (Levin & Beloborodov, 2003; Milosavljević & Loeb, 2004), through migration within a disk or a cluster (Levin, 2006; Fujii et al., 2010) to scatterings by massive perturbers and binary disruptions (Gould & Quillen, 2003; Ginsburg & Loeb, 2006; Perets et al., 2007; Generozov & Madigan, 2020).
In the context of TDEs, earlier works (e.g., Magorrian & Tremaine, 1999; Stone & Metzger, 2016) assumed an old stellar population and therefore truncated the stellar mass function at one solar mass. Here, we take a different approach. We assume that the stars in galactic centers have formed a relaxed cusp, and evaluate the resulting mass distribution of the tidally disrupted stars from such a cusp. Comparing this theory with inferred mass distributions from a large population of tidally disrupted stars, expected to be obtained through future observations, could provide insights into the degree of relaxation and mass segregation in galactic centers.
@@ 2 Steady-state Distributions
$R_{h}=\frac{GM_{BH}}{\sigma_{h}^{2}}\simeq 2\ {\rm pc}\left(\frac{M_{BH}}{M_{\rm MW}}\right)^{1/2}.$ (3)
The total enclosed stellar mass within the sphere of influence is roughly $M_{BH}$ (Binney & Tremaine, 1987; Merritt, 2004), leading to our simplified normalization $N_{\star}(1)\approx M_{BH}$.
2.1 The Present-Day Stellar Mass function
The present-day stellar mass function (PMF; Miller & Scalo, 1979; Kochanek, 2016) is given by
@@ 2.1 The Present-Day Stellar Mass function
where $\tau_{\rm gal}\sim 10\ \rm Gyr$ is the galaxy lifetime and $\tau_{\rm nuc}\sim 10m^{-2.5}\ {\rm Gyr}$ is the stellar nuclear timescale for a wide range of stellar masses. For stars with masses $m\gtrsim 40$, the stellar luminosities approach the Eddington limit (Sanyal et al., 2015) and their lifetimes become approximately constant, $\tau_{\rm nuc}(m\gtrsim 40)\sim{\rm Myr}$. We assume a Kroupa initial mass function (IMF; Kroupa, 2001). The broken power law of this IMF, as well as the distinction between stars with lifespans shorter and longer than the age of the galaxy, results in a broken power law PMF, ${\rm d}N_{\star}/{\rm d}m\propto m^{-\gamma}$, with
@@
Note that the short lifetimes of massive stars may prevent them from reaching a steady-state distribution, since their lifetimes can be shorter than the relaxation timescale. However, observations reveal a population of massive stars in the galactic center, where the local relaxation time exceeds their expected lifetimes (Ghez et al., 2003; Lu et al., 2006; Genzel et al., 2010). In the absence of a widely accepted model for how these stars arrived there, we adopt a simplified approach of assuming a relaxed stellar cusp, with the mass distribution described by the PMF. The finite lifetimes of stars are manifested in the steeper slope of the PMF for $m\gtrsim 1$ compared to the IMF (Eqs. 4 and 5).
2.2 BHs mass function

### v2

# Mass Segregation and Transient Formation in Nuclear Stellar Clusters

@@ Abstract
Supermassive black holes at the centers of galaxies occasionally disrupt stars or consume stellar-mass black holes (BHs) that wander too close, producing observable electromagnetic or gravitational wave signals. We examine how mass segregation impacts the rates and distributions of such events. Assuming a relaxed stellar cluster, composed of stars and stellar-mass BHs, we show that the tidal disruption rate of massive stars ($m\gtrsim{M_{\odot}}$) is enhanced relative to their abundance in the stellar population. For stars up to $m\approx 3{M_{\odot}}$, this enhancement is roughly $m/M_{\odot}$ and it is driven by segregation within the sphere of influence. Stars with masses $m\gtrsim 3{M_{\odot}}$, if relaxed, are predominantly scattered by more massive stellar-mass BHs, leading to a constant enhancement factor of $\approx 9$, independent of mass. This aligns with observational evidence suggesting an over-representation of massive stars in tidal disruption events. For stellar-mass BHs, we predict an enhancement factor scaling as $m_{\bullet}^{1/2}$ for plunges and $m_{\bullet}^{3/2}$ for extreme-mass-ratio inspirals (EMRIs). The power of one-half in both cases reflects the shorter relaxation times of heavier BHs, allowing them to segregate into the sphere of influence from greater distances, thereby increasing their abundance. The additional power in the EMRIs’ rate arises from the tendency of heavier BHs to circularize and sink inward more efficiently. Finally, we estimate the rate of main-sequence star inspirals and find that it favors low-mass stars ($m\lesssim{M_{\odot}}$). This seems compatible with the observationally estimated rate of quasiperiodic eruptions.
@@ 1 Introduction
The centers of galaxies harbor supermassive black holes (SMBHs) and their surrounding nuclear stellar clusters. These dense environments give rise to various electromagnetic and gravitational wave (GW) transients, such as tidal disruption events (TDEs; Rees, 1988; Gezari, 2021), quasiperiodic eruptions (QPEs; Miniutti et al., 2019; Arcodia et al., 2021), merging stellar-mass black hole (BH) binaries (Mapelli, 2021; Arca Sedda et al., 2023), and extreme-mass-ratio inspirals (EMRIs; Amaro-Seoane, 2018).
A star is tidally disrupted if it passes too close to the SMBH, namely closer than its tidal radius, where the SMBH’s tidal force and the star’s self-gravity are comparable (Hills, 1975)
@@
where $m$ and $R_{\star}$ are the mass and the radius of the star, respectively, and $R_{s}=2GM_{BH}/c^{2}$ is the SMBH Schwarzschild radius. We normalize $M_{BH}$ to the mass of Sgr A*, $M_{\rm MW}=4\times 10^{6}$ (Ghez et al., 2008; Gillessen et al., 2009), and use the main-sequence mass-radius relation, $R_{\star}\approx{R_{\odot}}(m/{M_{\odot}})^{0.8}$.
@@
Observationally, about $100$ TDEs have been discovered (Gezari, 2021; Hammerstein et al., 2023; Yao et al., 2023), primarily in the past decade, through wide-field surveys, such as the All-Sky Automated Survey for Supernovae (Jayasinghe et al., 2018) and the Zwicky Transient Facility (Bellm et al., 2019). The number of detected TDEs will increase significantly with the upcoming Vera Rubin Observatory Legacy Survey of Space and Time (Ivezic et al., 2019), which is expected to detect $\sim 10^{3}$ TDEs per year (van Velzen et al., 2011; Bricman & Gomboc, 2020).
@@
Furthermore, the recent discoveries of QPEs (Miniutti et al., 2019; Arcodia et al., 2021) uncovered a new class of transients from centers of galaxies, characterized by repeating X-ray emissions over periods of hours. The growing number of observed QPEs reveals the complexity and diversity of phenomena associated with these events (Arcodia et al., 2022; Miniutti et al., 2023; Miniutti, G. et al., 2023; Arcodia et al., 2024a; Chakraborty et al., 2024). Several formation models have been proposed, including mass transfer from white dwarfs (WDs; King, 2022) or stars (Metzger et al., 2022; Lu & Quataert, 2023; Linial & Sari, 2023), star-disk interactions (Xian et al., 2021; Franchini et al., 2023; Linial & Metzger, 2023), and disk instabilities (Raj & Nixon, 2021; Pan et al., 2022; Kaur et al., 2023). Evidence supporting the TDE–QPE connection, as suggested by Linial & Metzger (2023), has been observed (Chakraborty et al., 2021; Quintin et al., 2023; Bykov et al., 2024). Most recently, Nicholl et al. (2024) provided a direct link between these phenomena by detecting QPEs following a known TDE.
@@
In parallel, BHs accumulate in galactic centers and migrate inward due to dynamical friction, eventually merging with the central SMBH. These mergers produce GW signals in the mHz band, detectable by next-generation, space-based GW observatories, such as LISA (Amaro-Seoane et al., 2017, 2023) and TianQin (Luo et al., 2016). For these mergers, the minimal distance from the SMBH is $4R_{s}$, corresponding to the angular momentum of the mostly bound orbit (Hopman & Alexander, 2005), rather than the tidal radius associated with the disruption of stars. BHs that reach a smaller periapsis will rapidly plunge into the SMBH. In contrast, BHs on highly eccentric orbits with larger periapsis distance can dissipate energy via GW emission, circularize, and slowly descend toward the SMBH - such orbital evolution is known as EMRI (Hopman & Alexander, 2005; Amaro-Seoane, 2018).
The rates and characteristics of these transients depend on the distribution of stars and BHs around the SMBH, which has been extensively studied over the past half a century, since the pioneering works of Peebles (1972) and Bahcall & Wolf (1976), both theoretically (e.g., Alexander & Hopman, 2009; Linial & Sari, 2022) and numerically (e.g., Panamarev et al., 2019; Zhang & Amaro Seoane, 2025), as further described in Section 2.
The TDE rate, $10^{-4}-10^{-5}\ \rm yr^{-1}$ (Magorrian & Tremaine, 1999; Wang & Merritt, 2004; Holoien et al., 2015; Stone & Metzger, 2016; Kochanek, 2016; van Velzen, 2018; Stone et al., 2020; Bortolas et al., 2023; Yao et al., 2023), is determined by the replenishment of stars into highly eccentric orbits with periapsis $r_{p}\lesssim R_{t}$, mainly via two-body scatterings. However, other mechanisms can alter the TDE rate; for example, strong scatterings may reduce the TDE rate, given a steep profile density (Teboul et al., 2024; Kaur & Perets, 2024), while the presence of an SMBH binary can temporarily increase the TDE rate [^1] As well as the formation rate of EMRIs (see for example, Bode & Wegg, 2014; Naoz et al., 2022; Naoz & Haiman, 2023). (Chen et al., 2011; Li et al., 2015, 2019; Mockler et al., 2023; Melchor et al., 2024).
The population of disrupted stars is likely dominated by low-mass stars, with masses $m\lesssim{M_{\odot}}$, since they are more common and have a longer lifespan than more massive stars (Kochanek, 2016). Yet, there are observational indications of disruptions of more massive stars, with masses up to a few solar masses (e.g., Mockler et al., 2022; Hinkle et al., 2024; Wiseman et al., 2024).
@@
We highlight the impact of mass segregation on the rates and mass distribution of the different galacto-centric transients. For example, if the disrupted stars of all masses were to originate from roughly the radius of influence, their relative abundances would correspond to their prevalence in the stellar population. However, if mass segregation occurs, massive objects occupy more tightly bound orbits. The impact of such segregation is widely studied with regard to BHs and EMRI rates (e.g., Hopman & Alexander, 2006; Amaro-Seoane & Preto, 2011; Aharon & Perets, 2016; Raveh & Perets, 2021; Broggi et al., 2022; Rom et al., 2024). Analogously, mass segregation could increase the disruption rate of massive stars. But, unlike BHs, the relevance of segregation for massive stars is uncertain due to their short lifetimes.
@@
Nonetheless, massive stars are found in the center of our galaxy, despite the local relaxation time being longer than their expected ages. This is the known “paradox of youth” (Ghez et al., 2003; Lu et al., 2006; Genzel et al., 2010). Unfortunately, the mechanisms that enable massive stars to reside deep within the sphere of influence in our own galactic center are unclear. Various models have been suggested; including in situ formation (Levin & Beloborodov, 2003; Milosavljević & Loeb, 2004), migration within a disk or a cluster (Levin, 2006; Fujii et al., 2010), scattering by massive perturbers and binary disruptions (Gould & Quillen, 2003; Ginsburg & Loeb, 2006; Perets et al., 2007; Generozov & Madigan, 2020; Verberne et al., 2025; Generozov et al., 2025), and a recent massive BH binary merger (Akiba et al., 2024).
@@ 2 Steady-state Distributions
$R_{h}=\frac{GM_{BH}}{\sigma_{h}^{2}}\simeq 3\ {\rm pc}\left(\frac{M_{BH}}{M_{\rm MW}}\right)^{1/2},$ (2)
where we normalize $M_{BH}$ to the mass of Sgr A*, $M_{\rm MW}=4\times 10^{6}M_{\odot}$ (Ghez et al., 2008; Gillessen et al., 2009).
Under assumptions of spatial spherical symmetry, isotropic velocities, and weak two-body scattering dynamics, Bahcall & Wolf (1976) derived a steady-state “zero-flux” solution for a single-mass population. In this solution, the phase-space distribution is given by $f(E)\propto E^{p}$ with $p=1/4$, which corresponds to a number density $n(r)\propto r^{-\eta}$, with $\eta=3/2+p=7/4$ (hereafter, the BW cusp). This “zero-flux” solution is characterized by a vanishing particle flux and a constant energy flux (Shapiro & Lightman, 1976; Binney & Tremaine, 1987; Sari & Goldreich, 2006; Rom et al., 2023).
In realistic systems, where the energy (or semimajor axis) range is finite and an absorbing boundary condition is imposed near the SMBH, the steady-state solution has a finite, nonzero particle flux. Yet, this flux is significantly smaller - by roughly four orders of magnitude in a Milky Way-like galaxy - than the “naive” estimate, given by the number of bodies over the relaxation time (see Rom et al., 2023). This suppression reflects the ratio of energies in the scattering region, where two-body scattering dominates over GW emission (Amaro-Seoane, 2018; Sari & Fragione, 2019; Linial & Sari, 2023). Hence, the steady-state solution asymptotically approaches the “zero-flux” BW cusp.
In a following paper, Bahcall & Wolf (1977) generalized their calculation for multimass groups. Assuming that the most massive objects, $m_{\max}$, are the most abundant, a “zero-flux” solution is satisfied when the massive group follows the single-mass BW cusp, with $p_{\max}\approx 1/4$, while lighter objects, with mass $m_{i}<m_{\max}$, obtain shallower profiles (Bahcall & Wolf, 1977), satisfying
$\frac{p_{i}}{p_{\max}}\approx\frac{m_{i}}{m_{\max}}.$ (3)
However, low-mass stars are much more abundant than massive stars or compact objects; hence, they dominate the scattering and the massive objects sink toward the center of the cluster, due to dynamical friction, as encapsulated by Alexander & Hopman (2009) and Keshet et al. (2009) mass-segregated distributions. Linial & Sari (2022) revisited the impact of segregation in nuclear stellar clusters and derived a steady-state “zero-flux” solution, taking into account the dominance of different mass groups in different energy bins. This framework was adapted to a two-mass model by Rom et al. (2024), who used it to estimate the EMRI formation rate.
Furthermore, numerous numerical studies have examined the distributions around SMBHs, their evolution toward a steady state, and the consequent transients’ formation, using N Body simulations (e.g., Baumgardt et al., 2004a, b; Amaro-Seoane & Preto, 2011; Panamarev et al., 2019), Monte Carlo (MC) methods (e.g., Freitag & Benz, 2002; Freitag et al., 2006; Vasiliev, 2015; Bar-Or & Alexander, 2016; Balberg, 2024; Zhang & Amaro Seoane, 2024, 2025), or the Fokker-Planck (FP) formalism (e.g., Hopman & Alexander, 2006; O’Leary et al., 2009; Merritt, 2015a; Aharon & Perets, 2016; Vasiliev, 2017; Broggi et al., 2022). The formation of a BW cusp in a single-mass population surrounding an SMBH has been confirmed by multiple numerical studies since its original derivation (e.g., Freitag & Benz, 2002; Baumgardt et al., 2004a; Freitag et al., 2006; Vasiliev, 2015; Merritt, 2015b; Bar-Or & Alexander, 2016). In the multimass case, mass segregation of the heavier objects is generally observed, though the characteristics of the resulting distribution vary across the numerical studies, as further discussed in Section 2.4.
2.1 The Present-day Stellar-mass function
The PMF is given by (Miller & Scalo, 1979; Kochanek, 2016)
@@ 2.1 The Present-day Stellar-mass function
where $\tau_{\rm gal}\sim 10\ \rm Gyr$ is the galaxy lifetime and $\tau_{\rm nuc}\sim 10m^{-2.5}\ {\rm Gyr}$ is the stellar nuclear timescale over a wide range of stellar masses. For stars with masses $m\gtrsim 40$, the stellar luminosities approach the Eddington limit (Sanyal et al., 2015) and their lifetimes become approximately constant, $\tau_{\rm nuc}(m\gtrsim 40)\sim{\rm Myr}$. Given that the total enclosed stellar mass within the sphere of influence is roughly $M_{BH}$ (Binney & Tremaine, 1987; Merritt, 2004), we adopt a simplified normalization $N_{\star}(1)\approx M_{BH}$.
We assume a Kroupa IMF (Kroupa, 2001). The broken power law of this IMF, as well as the distinction between stars with lifespans shorter and longer than the age of the galaxy, results in a broken power-law PMF, ${\rm d}N_{\star}/{\rm d}m\propto m^{-\gamma}$, with
@@
Note that the short lifetimes of massive stars may prevent them from reaching a steady-state distribution, since their lifetimes can be shorter than the relaxation timescale. However, observations reveal a population of massive stars in the galactic center, where the local relaxation time exceeds their expected lifetimes (Ghez et al., 2003; Lu et al., 2006; Genzel et al., 2010). In the absence of a widely accepted model for how these stars arrived there, we adopt a simplified approach of assuming a relaxed stellar cusp, with the mass distribution described by the PMF. The finite lifetimes of stars are manifested in the steeper slope of the PMF for $m\gtrsim 1$ compared to the IMF (Equations 4 and 5).
2.2 BH mass function